The perimeter of a rectangle is 32 cm given that the length is 3times the width what is the area of rectangle

Answer by ikleyn(52786)

.

L + W = 32/2 = 16

L = 3W

==================>


3W + W = 16

4W     = 16

 W     = 16/4 = 4.


The width is 4 cm;  the length  is  3*4 = 12 cm.


The area = L*W = 12*4 = 48 cm^2.     ANSWER